# The Milky Way in context: The formation of galactic discs and chemical sequences from a cosmological perspective

### v1

@@ Abstract
We study the formation of chemical sequences in the stellar disc of Milky Way (MW)-mass galaxies in a full cosmological context with the Auriga simulations. We focus on the conditions giving rise to bi-modal $\alpha$-chemistry in the MW disc and the potential influence of mergers (e.g. Gaia-Enceladus, GSE). We find a wide diversity of chemical sequences, without correlation between the emergence of dichotomous $\alpha$-chemistry and GSE-like mergers. The transition between multiple $\alpha$-sequences is sequential, and is mediated by modulations in the star formation rate (SFR). In some cases, this can be caused by the starburst and subsequent quiescence induced by mergers. In others, by exhaustion or violent disruption of the gas disc. Realisations with singular sequences either lack significant modulations in their SFR, or form too late to have a significant high-$\alpha$ sequence. The metallicity overlap between the high-$\alpha$ and low-$\alpha$ sequences (as seen in the Solar neighbourhood of the MW) arises from accretion of metal-poor gas from the circum-galactic medium. This depends on gas disc thickness, with thinner discs losing their metal-poor extremities. Gas donation from singular gas-rich merger events are incapable of driving long-lived metal dilution ($\Delta\rm{[Fe/H]}\gtrsim 0.3$), and we rule-out this scenario for the low-$\alpha$ sequence in the MW. Finally, the shape of $\alpha$-sequences in the [Fe/H] versus [Mg/Fe] plane is related to long-term SFR trends, with a downward slanted locus (as is observed in the low-$\alpha$ sequence of the MW) owing to a sustained or declining SFR.
@@
pubyear: 2025pagerange: The Milky Way in context: The formation of galactic discs and chemical sequences from a cosmological perspectiveA
@@ 1 Introduction
Our vantage point within the Milky Way (MW) has allowed for comprehensive spectroscopic surveys into the chemo-dynamic properties of individual stars across large statistical samples and wide volumes (see RAVE Steinmetz et al. 2006, SEGUE & SEGUE-2 Yanny et al. 2009; Rockosi et al. 2022, Gaia-ESO Gilmore et al. 2012, LAMOST Deng et al. 2012, GALAH De Silva et al. 2015, APOGEE Majewski et al. 2017). This, combined with precise proper motions from Gaia (Gaia Collaboration et al., 2016; Lindegren et al., 2018; Gaia Collaboration et al., 2021), has led to a broader understanding of how the MW formed and established our Galaxy as a benchmark for studying the underlying physics of galaxy formation as a whole.
@@
High resolution spectroscopic studies of stars in the Solar neighbourhood have revealed that disc stars form a clear “bi-modal” distribution in the [Fe/H] versus [$\alpha$/Fe] chemical abundance planes, with two distinct sequences that are separated in their $\alpha$-chemistry but overlap in their metallicity distributions (e.g. Fuhrmann, 1998; Haywood et al., 2013; Anders et al., 2014; Nidever et al., 2014; Hayden et al., 2015). This bi-modality varies along with the location within the disc, both in terms of radius and height above or below the plane (e.g. see Queiroz et al., 2020). For inner radii (within $\approx 6\,$kpc), the sequences form a single track mediated by an under-populated gap (e.g. Hayden et al., 2015; Queiroz et al., 2020). For outer radii (beyond $\approx 12\,$kpc), the proportion of stars in the high-$\alpha$ sequence is strongly diminished (e.g. Bensby et al., 2011; Anders et al., 2014).
@@
The high-$\alpha$ and low-$\alpha$ sequences are well correlated with stars in the thick and thin discs respectively, and share similar qualitative properties. Whereas the high-$\alpha$ disc is uniformally old (having formed the majority of its stars earlier than 10 Gyr ago), centrally concentrated and vertically extended – the low-$\alpha$ disc is thin, radially extended and formed over a wide time period up to the present day (Haywood et al., 2013; Mackereth et al., 2017; Bonaca et al., 2020; Miglio et al., 2021; Queiroz et al., 2023). This is consistent with extra-Galactic observational evidence that galaxy formation proceeds in an inside-out and top-down manner (Franx et al., 2008; van Dokkum et al., 2013; Patel et al., 2013a; van der Wel et al., 2014; Lian & Luo, 2024), a result that is backed up by idealised and fully cosmological simulations (e.g. Bournaud et al., 2009; Brook et al., 2012; Bird et al., 2013). Yet, it is increasingly believed that the high/low-$\alpha$ and thick/thin discs are distinct populations of their own (e.g. see arguments in Robin et al., 1996; Jurić et al., 2008; Bovy et al., 2012, 2016). Nonetheless, the populations are often treated as interchangeable and chemical cuts are commonly employed to select the thick and thin disc populations.
@@
Whilst observational data shows that thick and thin discs are a frequent attribute of galaxies as a whole (e.g. Mould, 2005; Yoachim & Dalcanton, 2006; Tsukui et al., 2024), the same determination has not yet been made for chemical bi-modalities. Extra-galactic observations rely on integrated light, which blends stellar populations across disc regions (e.g. Scott et al., 2021). Furthermore, observations at higher redshift increasingly rely on gas-phase chemistry, which may not confer the chemical distribution of the stars. The one exception is our nearest neighbour M31. Early Keck/DEIMOS spectroscopy of 70 M31 stars found no sign of dichotomous chemistry (Escala et al., 2020), and near-infrared spectroscopy of around 100 stars from the JWST shows only a single sequence at high-$\alpha$ (Nidever et al., 2024). This is an early sign that dichotomous chemistry may not be an ubiquitous quality of massive disc galaxies, which suggests a diversity of formation channels.
@@
If the formation mechanism behind the chemical bi-modality was well known, one could debate whether it was a unique development to the MW. Despite many competing hypotheses, there has been no universal agreement on a dominant mechanism. These hypotheses fall into two categories: secular or episodic formation. An example of a secular mechanism would be the chemical evolution models of Schönrich & Binney (2009a, b). These find that low-metal and high-$\alpha$ stars form in the centrally-compact proto-Galaxy, and are then driven to the Solar neighbourhood by radial migration (as in Sellwood & Binney, 2002; Roškar et al., 2008). There, they overlap with younger, low-$\alpha$ stars in the radially extended thin disc. This scenario is backed up by chemical enrichment models in Sharma et al. (2021); Chen et al. (2023), $N$-body simulations in (Loebman et al., 2011), and could also explain other properties of the disc such as the radial dependence of the metallicity distribution function (Loebman et al., 2016). However, other models with radial migration fail to reproduce two chemical sequences (Minchev et al., 2013, 2014a; Johnson et al., 2021; Dubay et al., 2024).
@@
Otherwise, the two $\alpha$-sequences may form at the same time but in different Galactic environments. The authors of Clarke et al. (2019); Beraldo e Silva et al. (2021) argue that high-$\alpha$ stars could form in the dense intermediate-mass gas clumps that arise from disc fragmentation in the early Universe (see Elmegreen & Elmegreen, 2005; Elmegreen et al., 2007; Dekel et al., 2009; Livermore et al., 2012), within which the star formation rate (SFR) is raised and the $\alpha$-enrichment from core-collapse supernovae is enhanced. Then, low-$\alpha$ stars can form concurrently in a more evenly distributed gas disc. This scenario could simultaneously explain the elevated velocity dispersion of high-$\alpha$ stars via strong scattering processes (Bournaud et al., 2009; Amarante et al., 2020). However, observations from the CANDELS survey find that only up to 20 per cent of the integrated SFR comes from within clumps (see Wuyts et al. 2012, for galaxies over a similar redshift range and mass scale as the proto-MW).
@@
An episodic formation mechanism might suggest that a bi-modality requires a hiatus in the star formation between the formation of the high-$\alpha$ and low-$\alpha$ sequences (as in Haywood et al., 2016). This could be due either to a shrinking and then growth of the star-forming gas disc (i.e. Grand et al., 2018), or a suppression of the SFR following merger interactions or feedback from AGN (Beane, 2025; Beane et al., 2024). The pause in the star formation allows the delayed iron enrichment from Type-Ia supernovae to ‘catch up’, thereby reducing the ratio of [$\alpha$/Fe] prior to the onset of the low-$\alpha$ sequence. Indeed, there is tentative evidence for a quenching period around 8 Gyr ago from closed-box chemical evolution models in Haywood et al. (2018) and APOGEE DR17 data in Spitoni et al. (2024).
@@
It could also depend on distinct epochs of gas inflow, as in the two-infall model of (Chiappini et al., 1997, 2001). This proposes that the high-$\alpha$ sequence formed with a high SFR from a primordial rapid gas inflow, followed by a lower SFR from subsequent metal-poor gas accretion. The second infall ‘resets’ the ambient metallicity of the gas disc to lower values, thereby also explaining the signature [Fe/H] overlap between the two sequences in the Solar neighbourhood (see Grisoni et al., 2017; Spitoni et al., 2019, 2021, 2023). Prior works have discussed gas-rich mergers as the sources of a second gas infall (e.g. Richard et al., 2010; Snaith et al., 2016; Mackereth et al., 2018; Buck, 2020), or as a trigger that stimulates disc growth leading to low-$\alpha$ formation (see examples in Grand et al., 2018).
@@
The low-$\alpha$ disc may have formed from a prolonged fuelling from the circum-galactic medium (CGM). Studies have identified cold inflowing gas filaments (e.g. Anglés-Alcázar et al., 2017; Renaud et al., 2021), re-accretion of early gas outflows (e.g. Khoperskov et al., 2021), and smooth accretion from the CGM (e.g. Parul et al., 2025) as the dominant sources of gas during the formation of low-$\alpha$ sequences in simulations.
@@
Evidence from the stellar halo has now revealed the signature of an ancient massive merger, hereafter the Gaia-Sausage Enceladus or GSE (Helmi et al., 2018; Belokurov et al., 2018). This merger is predicted to have arrived in the MW around 8-11 Gyr ago (Vincenzo et al., 2019; Belokurov et al., 2020; Naidu et al., 2021; Xiang & Rix, 2022), roughly during the transition between the two $\alpha$-sequences (e.g. Linden et al., 2017; Sahlholdt et al., 2022), with a mass scale similar to that of the present-day Small Magellanic Cloud. Numerous works argue that the GSE played an important role in the history of the MW, with implications for its disc chemistry (e.g. Buck, 2020; Buck et al., 2023; Ciucă et al., 2024), the hot kinematics of the thick disc population (Helmi et al., 2018; Bonaca et al., 2020; Grand et al., 2020; Chandra et al., 2024) and the warp observed in the outer disc (Dillamore et al., 2022; Dodge et al., 2023; Deng et al., 2024). While evidence for the GSE merger is strong, its exact mass, infall angle, and impact on the proto-Galaxy remain uncertain. In particular, its exact relation to the chemical bi-modality is still circumstantial. This is a topic we intend to probe further.
@@ 2 Methods
The simulations analysed in this paper are drawn from the Auriga project (Grand et al., 2017, 2024), a suite of thirty magneto-hydrodynamic simulations of isolated MW-mass galaxies evolved within a full cosmological context. The initial conditions are generated at $z=127$ using the Gaussian white-noise realisation panphasia (Jenkins, 2013), and have a periodic cosmological box with a side length of 100100100\,100cMpc. The suite uses cosmological parameters from Planck Collaboration et al. (2014), which are $\Omega_{\rm m}=0.307$, $\Omega_{\rm b}=0.04825$, $\Omega_{\rm\Lambda}=0.693$ and a Hubble constant of $H_{0}=100h\,\text{km}^{-1}\,\text{Mpc}^{-1}$, where $h=0.6777$.
The Auriga simulations incorporate a comprehensive suite of sub-grid physics to model galaxy formation in a cosmological setting. These include treatments for a spatially uniform photoionizing UV background, primordial and metal-line cooling, star formation, stellar evolution, feedback from supernovae (SNe), supermassive black hole growth and feedback, as well as magnetohydrodynamics. A detailed overview of these physical prescriptions is provided in Grand et al. (2017, 2024). The simulations successfully reproduce a variety of key galactic properties expected in a cosmological context (Vogelsberger et al., 2013; Marinacci et al., 2014; Genel et al., 2014). The Auriga galaxies themselves exhibit characteristics consistent with observations of MW-like galaxies, including their halo mass–metallicity relations (Monachesi et al., 2019), vertical disc oscillations (Gómez et al., 2017), rotation curves and SFRs (Grand et al., 2017), and detailed chemodynamical structures in their central regions (Fragkoudi et al., 2020). Crucially, they also produce the presence of thick and thin disc components (Grand et al., 2018), both in terms of geometric, kinematic and chemical properties.
@@ 2.2 Post-processing
The group and subhalo properties are calculated using the subfind halo finder (Springel et al., 2001), and the virial masses and radii are derived using a sphere of mean density 200×200$\times$200 × the critical density of the universe. The evolution of halo structure is tracked across timesteps using the lhalotree merger tree algorithm (Springel et al., 2005). When necessary, inter-snapshot subalo properties are estimated using fitted splines. These are 3D cubic splines for subhalo orbits, and a PCHIP 1-D monotonic cubic interpolation for subhalo properties.
@@
The Auriga simulations track nine chemical species. We use Magnesium (Mg) for our estimation of $\alpha$-abundance because it has been shown to be a reliable tracer of the $\alpha$-process, with a high separation between the thick and thin discs (Jönsson et al., 2018). All chemical abundances in this work are normalised to Solar values using Asplund et al. (2009), with further corrections of order 0.4 dex used to lower the metallicity ratio ([Fe/H]). This work is often focused on the distribution of stars in the [Fe/H] versus [Mg/Fe] elemental abundance ratios, and for brevity we will hereafter refer to this simply as the ‘chemical abundance plane’.
@@
We focus exclusively on stars in the disc, following selection criteria based on both the stellar orbital circularity (using the definition for orbital circularity given in Abadi et al. 2003) and metallicity as described in the methods section of Orkney et al. (2023). Unless stated otherwise, we include stars from both in-situ and ex-situ sources, where the in/ex-situ designation is calculated based on whether the star particle is bound to the host halo at the earliest available snapshot. This is necessary to account for the presence of any ‘ex-situ disc’ (see Gómez et al., 2017). We note that the ex-situ contributions neither materially alter our reported results, nor do they create features in chemical abundance space that might be misinterpreted as independent chemical sequences (with Au-1 being the one exception). Furthermore, it is unlikely that the MW disc is dominated by ex-situ stars (e.g. Ruchti et al., 2011). The galaxies are aligned at every snapshot on the angular momentum of the youngest (formed in the last 3 Gyr) and innermost ($R_{\rm G}<0.1\times R_{200}$, where $R_{\rm G}$ is the galacto-centric radius) in-situ star particles.
@@
We calculate merger mass ratios based on the ratio between the total subfind masses (dark matter, gas and stars) as opposed to the commonly used ratio of $M_{200}$ virial masses. This is because the virial mass will often include massive substructures, and as a consequence the time-evolution of the mass ratio is more stochastic. The quoted merger mass ratio is then found as the maximum ratio after the satellite has joined the FoF group of the host structure. This usually occurs just after the satellite joins the FoF group, because gravitational tides and ram pressure stripping will begin to degrade it thereafter.

### v2

@@ Abstract
We study the formation of chemical sequences in stellar discs of Milky Way (MW)-mass galaxies in a full cosmological context with the auriga simulations, with implications for both the MW and external galaxies like M31. The analysis focuses on the conditions giving rise to bi-modal $\alpha$-chemistry and the potential influence of mergers (e.g. Gaia-Enceladus, GSE). We find a wide diversity of chemical sequences, without correlation between the emergence of dichotomous $\alpha$-chemistry and GSE-like mergers. The transition between multiple $\alpha$-sequences is sequential, and is mediated by modulations in the star formation rate (SFR). In some cases, this can be caused by the starburst and subsequent quiescence induced by mergers. In others, by exhaustion or violent disruption of the gas disc. Realisations with singular sequences either lack significant modulations in their SFR, or form too late to have a significant high-$\alpha$ sequence. The metallicity overlap between the high-$\alpha$ and low-$\alpha$ sequences (as seen in the Solar neighbourhood of the MW) arises from accretion of metal-poor gas from the circum-galactic medium. This depends on gas disc thickness, with thinner discs losing their metal-poor extremities. Gas donation from singular gas-rich merger events are incapable of driving long-lived metal dilution ($\Delta\rm{[Fe/H]}\gtrsim 0.3$), and we rule-out this scenario for the low-$\alpha$ sequence in the MW. Finally, the shape of $\alpha$-sequences in the [Fe/H] versus [Mg/Fe] plane is related to long-term SFR trends, with a downward slanted locus (as observed in the low-$\alpha$ sequence of the MW) owing to a sustained or declining SFR.
@@
pubyear: 2025pagerange: The Milky Way in context: The formation of galactic discs and chemical sequences from a cosmological perspectiveB
@@ 1 Introduction
Our vantage point within the Milky Way (MW) has allowed for comprehensive spectroscopic surveys into the chemo-dynamic properties of individual stars across large statistical samples and wide volumes (see RAVE rave, SEGUE & SEGUE-2 segue; segue-2, Gaia-ESO gaia-eso, LAMOST lamost, GALAH galah, APOGEE apogee). This, combined with precise proper motions from Gaia (gaia; gaiadr2; gaiadr3), has led to a broader understanding of how the MW formed and established our Galaxy as a benchmark for studying the underlying physics of galaxy formation as a whole.
@@
High resolution spectroscopic studies of stars in the Solar neighbourhood have revealed that disc stars form a clear “bi-modal” distribution in the [Fe/H] versus $\alpha$/Fe] chemical abundance planes, with two distinct sequences that are separated in their $\alpha$-chemistry but overlap in their metallicity distributions (e.g. fuhrmann1998; haywood2013; anders2014; nidever2014; hayden2015). This bi-modality varies along with the location within the disc, both in terms of radius and height above or below the plane (e.g. see queiroz2020). For inner radii (within $\approx 6\,$kpc), the sequences form a single track mediated by an under-populated gap (e.g. hayden2015; queiroz2020). For outer radii (beyond $\approx 12\,$kpc), the proportion of stars in the high-$\alpha$ sequence is strongly diminished (e.g. bensby2011; anders2014).
@@
The high-$\alpha$ and low-$\alpha$ sequences are well correlated with stars in the thick and thin discs respectively, and share similar qualitative properties. Whereas the high-$\alpha$ disc is uniformally old (having formed the majority of its stars earlier than 10 Gyr ago), centrally concentrated and vertically extended — the low-$\alpha$ disc is thin, radially extended and formed over a wide time period up to the present day (haywood2013; mackereth2017; bonaca2020; miglio2021; queiroz2023). This is consistent with extra-Galactic observational evidence that galaxy formation proceeds in an inside-out and top-down manner (franx2008; dokum2013; patel2013; vanderwel2014; lian2024), a result that is backed up by idealised and fully cosmological simulations (e.g. bournaud2009; brook2012; bird2013). Yet, it is increasingly believed that the high/low-$\alpha$ and thick/thin discs are distinct populations of their own (e.g. see arguments in robin1996; juric2008; bovy2012; bovy2016). Nonetheless, the populations are often treated as interchangeable and chemical cuts are commonly employed to select the thick and thin disc populations.
@@
Whilst observational data shows that thick and thin discs are a frequent attribute of galaxies as a whole (e.g. mould2005; yoachim2006; tsukui2024), the same determination has not yet been made for chemical bi-modalities. Extra-galactic observations rely on integrated light, which blends stellar populations across disc regions (e.g. scott2021). Furthermore, observations at higher redshift increasingly rely on gas-phase chemistry, which may not confer the chemical distribution of the stars. The one exception is our nearest neighbour M31. Early Keck/DEIMOS (Deep Imaging Multi-Object Spectrograph) spectroscopy of 70 M31 stars found no sign of dichotomous chemistry (escala2020), and near-infrared spectroscopy of around 100 stars from the James Webb Space Telescope (JWST) shows only a single sequence at high-$\alpha$ (nidever2024). This is an early sign that dichotomous chemistry may not be an ubiquitous quality of massive disc galaxies, which suggests a diversity of formation channels.
@@
If the formation mechanism behind the chemical bi-modality was well known, one could debate whether it was a unique development to the MW. Despite many competing hypotheses, there has been no universal agreement on a dominant mechanism. These hypotheses fall into two categories: secular or episodic formation. An example of a secular mechanism would be the chemical evolution models of schonrich2009a; schonrich2009. These find that low-metal and high-$\alpha$ stars form in the centrally-compact proto-Galaxy  [^1] In which a proto-galaxy is defined as the most ancient progenitor to modern-day galaxies, that has undergone minimal star formation or chemical enrichment (see for e.g. belokurov2022), and are then driven to the Solar neighbourhood by radial migration (as in sellwood2002; rovskar2008). There, they overlap with younger, low-$\alpha$ stars in the radially extended thin disc. This scenario is backed up by chemical enrichment models in sharma2021; chen2023, $N$-body simulations in (loebman2011), and could also explain other properties of the disc such as the radial dependence of the metallicity distribution function (loebman2016). However, other models with radial migration fail to reproduce two chemical sequences (minchev2013; minchev2014; johnson2021; dubay2024).
@@
Otherwise, the two $\alpha$-sequences may form at the same time but in different Galactic environments. The authors of clarke2019; beraldo2021 argue that high-$\alpha$ stars could form in the dense intermediate-mass gas clumps that arise from disc fragmentation in the early Universe (see elmegreen2005; elmegreen2007; dekel2009; livermore2012), within which the star formation rate (SFR) is raised and the $\alpha$-enrichment from core-collapse supernovae is enhanced. Then, low-$\alpha$ stars can form concurrently in a more evenly distributed gas disc. This scenario could simultaneously explain the elevated velocity dispersion of high-$\alpha$ stars via strong scattering processes (bournaud2009; amarante2020b). However, observations from the CANDELS survey (The Cosmic Assembly Near-infrared Deep Extragalactic Legacy Survey) find that only up to 20 per cent of the integrated SFR comes from within clumps (see wuyts2012, for galaxies over a similar redshift range and mass scale as the proto-MW).
@@
An episodic formation mechanism might suggest that a bi-modality requires a hiatus in the star formation between the formation of the high-$\alpha$ and low-$\alpha$ sequences (as in haywood2016). This could be due either to a shrinking and then growth of the star-forming gas disc (i.e. grand2018), or a suppression of the SFR following merger interactions or feedback from AGN (beane2024a; beane2024b). The pause in the star formation allows the delayed iron enrichment from Type-Ia supernovae to ‘catch up’, thereby reducing the ratio of $\alpha$/Fe] prior to the onset of the low-$\alpha$ sequence. Indeed, there is tentative evidence for a quenching period around 8 Gyr ago from closed-box chemical evolution models in haywood2018 and APOGEE DR17 data in spitoni2024.
@@
It could also depend on distinct epochs of gas inflow, as in the two-infall model of (chiappini1997; chiappini2001). This proposes that the high-$\alpha$ sequence formed with a high SFR from a primordial rapid gas inflow, followed by a lower SFR from subsequent metal-poor gas accretion. The second infall ‘resets’ the ambient metallicity of the gas disc to lower values, thereby also explaining the signature [Fe/H] overlap between the two sequences in the Solar neighbourhood (see grisoni2017; spitoni2019; spitoni2021; spitoni2023). Prior works have discussed gas-rich mergers as the sources of a second gas infall (e.g. richard2010; snaith2016; mackereth2018; buck2020), or as a trigger that stimulates disc growth leading to low-$\alpha$ formation (see examples in grand2018).
@@
The low-$\alpha$ disc may have formed from a prolonged fuelling from the circum-galactic medium (CGM). Studies have identified cold inflowing gas filaments (e.g. angles2017; agertz2021b), re-accretion of early gas outflows (e.g. khoperskov2021), and smooth accretion from the CGM (e.g. parul2025) as the dominant sources of gas during the formation of low-$\alpha$ sequences in simulations.
@@
Evidence from the stellar halo has now revealed the signature of an ancient massive merger, hereafter the Gaia-Sausage Enceladus or GSE (helmi2018; belokurov2018). This merger is predicted to have arrived in the MW around 8-11 Gyr ago (vincezo2019; belokurov2020; naidu2021; xiang2022), roughly during the transition between the two $\alpha$-sequences (e.g. linden2017; sahlholdt2022), with a mass scale similar to that of the present-day Small Magellanic Cloud. Numerous works argue that the GSE played an important role in the history of the MW, with implications for its disc chemistry (e.g. buck2020; buck2023; ciuca2024), the hot kinematics of the thick disc population (helmi2018; bonaca2020; grand2020; chandra2024) and the warp observed in the outer disc (dillamore2022; dodge2023; deng2024). While evidence for the GSE merger is strong, its exact mass, infall angle, and impact on the proto-Galaxy remain uncertain. In particular, its exact relation to the chemical bi-modality is still circumstantial. This is a topic we intend to probe further.
@@ 2 Methods
The simulations analysed in this paper are drawn from the auriga project (Auriga; grand2024), a suite of thirty magneto-hydrodynamic simulations of isolated MW-mass galaxies evolved within a full cosmological context. The initial conditions are generated at $z=127$ using the Gaussian white-noise realisation panphasia (Panphasia), and have a periodic cosmological box with a side length of 100100\,cMpc. The suite uses cosmological parameters from Planck2014, which are $\Omega_{\rm m}=0.307$, $\Omega_{\rm b}=0.04825$, $\Omega_{\rm\Lambda}=0.693$ and a Hubble constant of $H_{0}=100h\,\text{km}^{-1}\,\text{Mpc}^{-1}$, where $h=0.6777$.
The auriga simulations incorporate a comprehensive suite of sub-grid physics to model galaxy formation in a cosmological setting. These include treatments for a spatially uniform photoionizing UV background, primordial and metal-line cooling, star formation, stellar evolution, feedback from supernovae (SNe), supermassive black hole growth and feedback, as well as magnetohydrodynamics. A detailed overview of these physical prescriptions is provided in Auriga; grand2024. The simulations successfully reproduce a variety of key galactic properties expected in a cosmological context (vogelsberger2013; marinacci2014; genel2014). The auriga galaxies themselves exhibit characteristics consistent with observations of MW-like galaxies, including their halo mass–metallicity relations (monachesi2019), vertical disc oscillations (gomez2017), rotation curves and SFRs (Auriga), and detailed chemodynamical structures in their central regions (fragkoudi2020). Crucially, they also produce the presence of thick and thin disc components (grand2018), both in terms of geometric, kinematic and chemical properties.
The auriga galaxy formation models employ a two-phase interstellar medium, consisting of cold clouds immersed in a hot gaseous background. Star formation occurs stochastically in gas above a threshold density of $0.13\,m_{\rm p}\,\mathrm{cm}^{-3}$, with the target stellar particle mass matching the initial gas cell resolution. Each star particle represents a simple stellar population following a Chabrier initial mass function (chabrier2003). Mass and metals are returned to the gas via sub-grid models for AGB winds, Type-II, and Type-Ia supernovae. Here, the returns from Type II supernovae are assumed to occur instantaneously, whilst those from Type Ia supernovae follow a time-delay distribution resulting in a characteristic enrichment timescale of $\approx 1\,$Gyr. Further details on the physics prescriptions, including the adopted yield tables, are provided in the auriga public release (grand2024) and introductory paper (Auriga). grand2024 also discusses how these choices compare to the physics implemented in the TNG simulations (nelson2019).
@@ 2.2 Post-processing
The group and subhalo properties are calculated using the subfind halo finder (subfind), and the virial masses and radii are derived using a sphere of mean density $200\times$ the critical density of the universe. The evolution of halo structure is tracked across timesteps using the lhalotree merger tree algorithm (springel2005). When necessary, inter-snapshot subalo properties are estimated using fitted splines. These are 3-D cubic splines for subhalo orbits, and a PCHIP (Piecewise Cubic Hermite Interpolating Polynomial) 1-D monotonic cubic interpolation for subhalo properties.
@@
The auriga simulations track nine chemical species. We use Magnesium (Mg) for our estimation of $\alpha$-abundance because it has been shown to be a reliable tracer of the $\alpha$-process, with a high separation between the thick and thin discs (jonsson2018). All chemical abundances in this work are normalised to Solar values using asplund2009, with further corrections of order 0.4 dex used to lower the metallicity ratio ([Fe/H]). This work is often focused on the distribution of stars in the [Fe/H] versus [Mg/Fe] elemental abundance ratios, and for brevity we will hereafter refer to this simply as the ‘chemical abundance plane’.
We focus exclusively on stars in the disc, following selection criteria based on both the stellar orbital circularity (using the definition for orbital circularity given in abadi2003) and metallicity as described in the methods section of orkney2023. To summarise, we calculate the circularity distribution of all star particles and assume all retrograde stars belong to the stellar halo. Then, we mirror this distribution around $\eta=0$ to find the probability that prograde stars belong to the halo. Prograde star particles are probabilistically assigned to the halo using a random number generator, and the remaining stars are classified as disc. Additionally, we enforce that the retrograde and prograde halo components share an equivalent metallicity distribution function, though we note this has little implication on the final selections.
@@
Unless stated otherwise, we include stars from both in-situ and ex-situ sources, where the in/ex-situ designation is calculated based on whether the star particle is bound to the host halo at the earliest available snapshot. This is necessary to account for the presence of any ‘ex-situ disc’ (see gomez2017). We note that the ex-situ contributions neither materially alter our reported results, nor do they create features in chemical abundance space that might be misinterpreted as independent chemical sequences (with Au-1 being the one exception). Furthermore, it is unlikely that the MW disc is dominated by ex-situ stars (e.g. ruchti2011). The galaxies are aligned at every snapshot on the angular momentum of the youngest (formed in the last 3 Gyr) and innermost ($R_{\rm G}<0.1\times R_{200}$, where $R_{\rm G}$ is the galacto-centric radius) in-situ star particles.
@@
We calculate merger mass ratios based on the ratio between the total subfind masses (dark matter, gas and stars) as opposed to the commonly used ratio of $M_{200}$ virial masses. This is because the virial mass will often include massive substructures, and as a consequence the time-evolution of the mass ratio is more stochastic. The quoted merger mass ratio is then found as the maximum ratio after the satellite has joined the FoF (Friends of Friends) group of the host structure. This usually occurs just after the satellite joins the FoF group, because gravitational tides and ram pressure stripping will begin to degrade it thereafter.
[…]